# PoU: Proof-of-Use to Counter Tool-Call Hacking in DeepResearch Agents

### v1

@@ 3.1.3. Reasoning Process
At the beginning of each iteration, the agent firstly performs reasoning enclosed within a <think> tag. For the first step, the model generates initial planning and reasoning solely based on the user question. From the second step onward, each <think> block must begin with a explicit verdict and citation declaration, which is treated as a unified contract:
@@ 3.2. Proof-of-use Rewards
3.2.1. Unified Step Contract for Citation Reward
According to the unified contract (denoted in 3.1.3) of each step, all proxy responses return dictionaries containing a stable id, enabling deterministic auditing. We define the Cite Reward $Reward_{cite}^{t}$ in step $t$:
(1) $\begin{split}Reward_{cite}^{t}=\begin{cases}+1,&\text{if }\texttt{parse\_ok}\land\texttt{consistency\_ok}\land\texttt{ids\_valid},\\ -1,&\text{otherwise.}\end{cases}\end{split}$
@@ 3.2.1. Unified Step Contract for Citation Reward
$Reward_{cite}=\begin{cases}\frac{\sum_{t=2}^{T}Reward_{cite}^{t}}{T-1},&T>1,\\[4.0pt] 0,&T=1,\end{cases}$
@@
3.3. Reference-Linked Evidence Perturbations
To evaluate whether the model’s decision truly depends on the cited evidence rather than spurious features, we design a perturbation-based reward that measures the directional sensitivity of the model’s helpfulness prediction $p_{\theta}(\texttt{yes})$ to evidence degradation:
@@ 3.3. Reference-Linked Evidence Perturbations
Specifically, for the YES case, we degrade the cited evidence set $S_{t}$ by replacing their contents with topic-unrelated snippets, and observe the probability change $\Delta q=q^{\prime}-q$, where $q=p_{\theta}(\texttt{yes}\mid R_{\mathrm{real}})$ and $q^{\prime}=p_{\theta}(\texttt{yes}\mid R_{\mathrm{pert}})$. A model that truly relies on the cited evidence should decrease its helpfulness confidence ($\Delta q<0$) after the evidence is corrupted.
For the complementary NO case, we apply the opposite perturbation: when the model initially predicts the helpfulness token as no, we randomly select one document from the proxy-returned list and replace its content with an LLM-generated snippet that is semantically relevant to the query but not necessarily factual. This synthetic “semantic lure” mimics a plausible piece of supportive evidence.
Since each rollout may contain multiple steps, the perturbation reward is computed under a fixed budget $B$:
@@
and we randomly sample $B$ steps from those with $Reward_{cite}=1$ (from step 2 to $T$) to apply perturbation. We set $B_{\max}=1$ for main experiments (thus $B=1$), and study $B=2$ in ablations to assess stability–cost trade-offs.
The final rollout-level perturbation reward is the mean of the selected $B$ results.
3.4. Answer–Citation Alignment
The large action space and delayed rewards make the agent struggle to assign credit properly, often resorting to shortcut guesses rather than grounded reasoning. To overcome the sparse credit assignment and shortcut reasoning issues, PoU directly ties the final reward to the causal consistency between the produced answer and the cited evidence through an Answer–Citation Alignment objective.
We compute the answer–evidence alignment reward as:
(3) $\begin{split}Reward_{ac}=\begin{cases}0,&\text{if }\sum_{t}\mathbb{I}[Reward_{cite}^{(t)}{=}1]=0\\ &\text{or }T{=}1,\\[3.0pt] \mathcal{J}(q,\hat{a},\mathcal{E})\in\{0,0.5,1\},&\text{otherwise.}\end{cases}\end{split}$
where $\mathcal{J}$ denotes the external LLM judge that scores whether $\hat{a}$ can be reasonably derived from the aggregated evidence set $\mathcal{E}$ of <helpful>yes</helpful> steps.
The final answer-level reward combines factual adequacy and correctness as
(4) $Reward_{a}=\frac{Reward_{ac}+Reward_{ans}}{2},$
where $Reward_{ans}$ is the F1 score between the predicted and gold answers.
We avoid overlap-based metrics (BLEU, ROUGE) for $Reward_{ac}$ since they can be trivially optimized by surface copying rather than genuine reasoning.
@@ 3.4.1. Overall Objective
(5) $Reward_{\mathrm{final}}=\begin{cases}\frac{Reward_{cite}+Reward_{pt}+Reward_{a}}{3},&\text{if format is valid,}\\ -1,&\text{otherwise.}\end{cases}$
@@
All components are jointly optimized using Group Relative Policy Optimization (GRPO)(Shao et al., 2024). Unlike conventional agentic RL pipelines that fully mask tool outputs during training, our rollout retains the actually cited evidence snippets (those referenced via <ref>) unmasked. This aims to encourage the model to directly condition on verified causal dependency between retrieved evidence and reasoning.
@@ 3.5. Startup Fine-Tuning
To initialize a reliable agent under our PoU reasoning framework, we generate expert trajectories by a strong LLM (GPT-5). These multi-step reasoning trajectories are reformatted into our standardized interaction schema, where each turn explicitly follows the PoU contract.
@@
The resulting dataset thus consists of verifiable, well-grounded reasoning episodes that adhere to our explicit stepwise contract. We fine-tune the base model on this curated corpus to cold-start the PoU agent.
@@ 4.2. Baselines
We benchmark PoU against a diverse set of recent deep-research agents, spanning both retrieval–reasoning paradigms and RL-based search models.
(1) Retrieval–Reasoning RAG Baselines.
Routing RAG adopts a tool-routing design, where a controller dynamically selects the appropriate retriever (e.g., local corpus, web, or KG) based on query type and reasoning context. All-in-One RAG, by contrast, fuses multiple retrievers into a unified interface, jointly encoding multi-source evidence for direct reasoning without explicit routing. Both follow the ReAct (Yao et al., 2023) framework with three reasoning–retrieval turns.
@@
Search-o1 (Li et al., 2025a) performs multi-hop search by iteratively generating sub-queries and retrieving concise snippets rather than full documents. Search-R1 (Jin et al., 2025) employs reinforcement learning for question answering, leveraging a Wikipedia retriever; it has both base and instruct variants. R1-Searcher (Song et al., 2025) also uses RL, but restricts retrieval to Wikipedia via site-specific filters, followed by multi-page summarization.
(3) Multi-Tool Deep-Research Agents.
DeepResearcher (Zheng et al., 2025) and Chain-of-Agent (CoA) (Li et al., 2025b) represent the latest generation of multi-source deep-research agents, operating in heterogeneous tool environments. PoU builds upon this line, introducing explicit proof-of-usefulness supervision and unified reasoning contracts to ensure verifiable evidence grounding.
4.3. Baselines
To evaluate the effectiveness of PoU, we benchmark it against a range of recent deep-research agents. Routing RAG and All-in-One RAG represent two complementary retrieval-reasoning paradigms. Routing RAG adopts a tool-routing design, where the controller dynamically decides which retriever (e.g., local corpus, web search, or knowledge graph) to invoke at each step according to the query type and reasoning context. This enables flexible retrieval planning but still relies on pre-defined routing heuristics. All-in-One RAG, in contrast, fuses all retrievers into a unified interface that jointly encodes multi-source evidence, letting the model reason directly over concatenated contexts without explicit routing. All RAG baseline following ReAct(Yao et al., 2023) with 3 turns, reflecting a fully integrated retrieval-generation pipeline.
Beyond these two RAG baselines, Search-o1(Li et al., 2025a) follows a multi-hop reasoning process where the model sequentially generates intermediate queries and retrieves short evidence snippets rather than full documents. Search-R1(Jin et al., 2025) employs reinforcement learning for question answering, using a retriever to gather Wikipedia-based evidence, with base and instruct variants depending on the initialization of the actor model. R1-Searcher(Song et al., 2025) also applies RL but confines the search scope to Wikipedia by appending site-specific constraints and summarizing multiple retrieved pages. Finally, DeepResearcher(Zheng et al., 2025) and Chain-of-Agent (CoA)(Li et al., 2025b) are two most recent deepresearch baselines with multi-source tool environments.
@@ 4.4. Implementation Details
We use GPT-4o-mini as the external judge model for evaluating answer–evidence alignment and factual consistency. All other agentic modules (browser, KG proxy, perturbations LLM) are implemented with Qwen-3-8B, deployed via vLLM inference servers. The training backbone model is Qwen-2.5-7B-Instruct. Prior to reinforcement learning, supervised fine-tuning is performed with a batch size = 256 for 3 epochs. For reinforcement learning, we adopt VERL with the GRPO algorithm. Each update uses 8 parallel rollouts, distributed across 32 × A100 GPUs. The $Reward_{pt}$ budget $B$ is 1. The maximum reasoning step is set to 10, and each retrieval call returns the top-5 documents or chunks per query. The KG proxy retrieves one-hop subgraph information from Wikidata and WikiPedia by ToG-2(Ma et al., 2025) system, providing semantically condensed neighborhood summaries as context. In the OODT experiments, we additionally introduce a local biochemical Proxy, which is built from a biochemical corpus derived from PubMed and serves as a local retrieval interface, serving as the fixed local knowledge base for domain-specific retrieval. The local_proxy is turn-off for non-medical domain tasks.
@@ 4.5. Main Results
Across all seven datasets—including both in-domain (HotpotQA, 2Wiki) and out-of-domain (NQ, TQ, MuSiQue, Bamboogle, PopQA) benchmarks—our proposed Proof-of-Usefulness (PoU) agent consistently outperforms prior RAG and RL-trained baselines. This advantage is not limited to the in-domain subset: on the out-of-domain benchmarks, PoU achieves clear gains in both factual F1 and LLM-as-judge (LM) metrics, demonstrating superior generalization beyond its training distribution.
Interestingly, even on datasets such as NQ and TQ—which serve as in-domain data for the DeepResearcher baseline—our model achieves higher overall scores. This observation suggests that PoU’s improvement does not rely on memorizing domain-specific knowledge, but rather stems from stronger evidence–reasoning alignment. The model effectively learns to utilize retrieved evidence and to generalize its reasoning strategies to novel domains, indicating that PoU captures transferable reasoning competence rather than dataset-specific experience.
We also observe that All-in-One RAG consistently outperforms Routing RAG across all evaluation sets. This finding corroborates the observation in RAG in the Wild that naïve or heuristic routing—without explicit optimization or supervision—tends to be inefficient or even counterproductive. Although routing appears intuitively beneficial by narrowing the retrieval scope and reducing search cost, it also risks prematurely discarding relevant information.
In our context, the PoU agent operates in a multi-source environment where routing becomes an additional latent variable in the decision space. Compared with traditional deep-research agents, this introduces a new layer of complexity: each reasoning step must now decide not only what to retrieve but also where to retrieve from. Such an expanded action space grows exponentially with the number of information sources, leading to inefficient exploration when the routing policy is not explicitly optimized—precisely the phenomenon we aim to address through structured reasoning and contract-based training under the PoU framework. By explicitly encoding citation validity and reasoning consistency in every decision step, PoU regularizes the search process and prevents the agent from diverging in the enlarged action space. This design enables the agent to leverage multi-source evidence effectively without suffering from the instability and inefficiency commonly observed in naive routing-based systems, ultimately yielding superior performance and generalization ability in multi-source environments.
@@ 4.6.1. Rewards Ablation
Figure 1 presents the ablation study of PoU under different reward configurations, where each reward component is selectively removed or adjusted. We observe that increasing the perturbation budget to B=2B{=}2 leads to a more stable training trajectory and a higher final performance. This confirms the effectiveness of the perturbation-based reward $Reward_{pt}$: by actively perturbing the cited evidence, the model learns to rely on genuinely supportive information rather than superficially acknowledging tool responses. Such perturbation helps prevent the agent from “pretending” to use tool feedback, thereby maintaining a healthy and optimizable training regime instead of exploiting spurious reward shortcuts. However, since the rollout cost of GRPO scales approximately multiplicatively with the perturbation budget, we adopt B=1B{=}1 as a balanced trade-off between computational efficiency and stability.
Furthermore, removing the answer–citation alignment term $Reward_{ac}$ results in unstable training and a sudden collapse in later stages. This indicates that linking the final answer with intermediate evidence provides an essential smoothing and grounding signal for the policy. Without this alignment, the model easily drifts away from verifiable reasoning behaviors, demonstrating the critical role of answer–evidence coupling in maintaining robust optimization.
Figure 1. Ablation study of PoU. Comparison under different reward configurations: default perturbation budget B=1B{=}1, extended B=2B{=}2, and variants removing $Reward_{pt}$ or $Reward_{ac}$. Increasing the perturbation budget to B=2B{=}2 yields more stable convergence and a higher upper bound, validating the effectiveness of the perturbation-based reward $Reward_{pt}$. Perturbation prevents the agent from “pretending” to read tool feedback and keeps optimization in a healthy, learnable regime rather than exploiting spurious shortcuts. However, since GRPO rollout cost scales multiplicatively with $B$, we adopt B=1B{=}1 as a practical trade-off. Removing the answer–citation alignment term $Reward_{ac}$ causes the model to collapse in later stages, confirming that explicit evidence–answer alignment provides a crucial smoothing and grounding signal for stable training.
4.6.2. Out-of-domain (OODT) Experiments
To further validate the generalization ability of our approach in vary environments, we conducted additional experiments under an out-of-domain (OOD+OODT) setting. Specifically, we introduced a fourth proxy—a biomedical local search proxy—which was never seen during training. We then compared our model with existing baselines on two biomedical QA datasets, BioASQ(Krithara et al., 2023) and PQArefEval(Bašaragin et al., 2024). This setup constitutes an OOD+OODT scenario (out-of-domain data and out-of-domain tools), enabling us to assess how well the model generalizes when both the data distribution and the available tool set differ from those encountered during training.
Dataset Zero-shot RAG DeepResearcher+ DeepResearcher PoU (Ours)
BioASQ 13.5 25.8 37.1 38.9 45.6
PQArefEval 18.4 31.5 42.4 42.1 48.7
Table 3. OOD+OODT evaluation on biomedical QA datasets (BioASQ, PQArefEval). The experiments introduce a fourth proxy, a biomedical local search proxy, to test generalization across both domain and toolset shifts.
As shown in Table 3, even though DeepResearcher+ was trained in a more complex 3-tool environment, its performance is notably inferior to its original 2-tool counterpart (DeepResearcher) on both biomedical datasets. In contrast, our proposed PoU model achieves the best results by a clear margin, demonstrating strong domain-level and tool-level generalization. This finding highlights an important phenomenon: as the number of available tools increases, the action space of the agent expands combinatorially. Without explicit process-level reward signals, such enlarged search space can easily lead to policy collapse or unstable reasoning behavior. In contrast, the PoU training paradigm—through explicit evidence-reasoning alignment and structured reward design—successfully stabilizes the learning process and preserves robustness under tool and domain distribution shifts, indicating the Tool-call Hacking is significantly mitigated.

### v2

@@ 3.2.3. Reasoning Process
At the beginning of each iteration, the agent firstly performs reasoning enclosed within a <think> tag. For the first step, the model generates initial planning and reasoning solely based on the user question. From the second step onward, each <think> block must begin with a explicit helpfulness verdict and citation declaration, which is treated as a unified protocol:
@@ 3.3. Proof-of-Use Rewards
3.3.1. Unified Step protocol for Citation Reward
According to the unified protocol (denoted in 3.2.3) of each step, all proxy responses return dictionaries containing an unique ID, enabling deterministic auditing. We define the Cite Reward $R_{cite}^{t}$ in step $t$:
(2) $\begin{split}R_{cite}^{t}=\begin{cases}+1,&\text{if }\texttt{parse\_ok}\land\texttt{consistency\_ok}\land\texttt{ids\_valid},\\ -1,&\text{otherwise.}\end{cases}\end{split}$
@@ 3.3.1. Unified Step protocol for Citation Reward
$R_{cite}=\begin{cases}\frac{\sum_{t=2}^{T}R_{cite}^{t}}{T-1},&T>1,\\[4.0pt] 0,&T=1,\end{cases}$
@@
3.3.2. Evidence-Sensitive Reward
To evaluate whether the model’s decision truly depends on the cited evidence rather than spurious features, we design a perturbation-based reward that measures the sensitivity of the model’s helpfulness prediction $p_{\theta}(\texttt{yes})$ to evidence change at step $t$:
@@ 3.3.2. Evidence-Sensitive Reward
Specifically, for the YES case, we degrade the cited evidence set $S_{t}$ by replacing their contents with topic-unrelated snippets, and observe the probability change $\Delta q=q^{\prime}-q$, where $q=p_{\theta}(\texttt{yes}\mid rsp_{\mathrm{real}})$, $q^{\prime}=p_{\theta}(\texttt{yes}\mid rsp_{\mathrm{pert}})$, $rsp_{real}$ is the real tool responses and $res_{pert}$ is the real response with partial replacement. A model that truly relies on the cited evidence should decrease its helpfulness confidence ($\Delta q<0$) after the evidence is corrupted.
For the complementary NO case, we apply the opposite perturbation: when the model initially predicts the helpfulness token as no, we randomly select one document from the proxy-returned list and replace its content with an LLM-generated snippet that is semantically relevant to the query but not necessarily factual. This synthetic “semantic lure” mimics a plausible piece of supportive evidence. A model that truly relies on the cited evidence should increase its helpfulness confidence ($\Delta q>0$) after lure evidence are inserted.
Since each rollout may contain multiple steps, the Evidence-Sensitive reward is computed under a fixed budget gate $B$:
@@
where we randomly sample $B$ steps from those with $R_{cite}=1$ (from step 2 to $T$) to apply perturbation. We set $B_{\max}=1$ for main experiments (thus $B=1$), and study $B=2$ in ablations to assess stability–cost trade-offs.
The final rollout-level Evidence-Sensitive reward is defined as the mean over the selected perturbation outcomes:
(4) $R_{pt}=\frac{1}{B}\sum_{t\in\mathcal{S}_{B}}R_{pt}^{(t)}.$
In practice, by caching prefix KV states and original <helpful> logits, evidence perturbation is reduced to a tail-only forward pass; see Appendix LABEL:app:useful_cache.
3.3.3. Answer-Support Alignment Reward
We consider two complementary training signals: an Answer-Support Alignment reward $R_{ans}$ and a common Answer-Correctness reward $R_{anc}$.
The large action space and delayed rewards make the agent struggle to assign credit properly, often resorting to shortcut guesses rather than grounded reasoning. To overcome the sparse credit assignment and shortcut reasoning issues, PoU directly provides dense and smooth supervision by encouraging answers grounded in retrieved evidence through the Answer-Support alignment objective.
We compute the Answer–Support reward as:
(5) $\begin{split}R_{ans}=\begin{cases}0,&\text{if }\sum_{t}\mathbb{I}[R_{cite}^{(t)}{=}1]=0\\ &\text{or }T{=}1,\\[3.0pt] \mathcal{J}(q,\hat{a},\mathcal{E})\in[0,1],&\text{otherwise.}\end{cases}\end{split}$
where $\mathcal{J}$ denotes the external judge model measures the Answer-Support Alignment score of the generated answer $\hat{a}$ based on the cited evidence $\mathcal{E}$.
Here, we opt to the Faithfulness [^1] https://docs.ragas.io/en/stable/concepts/metrics/available_metrics/faithfulness/ metric of Ragas(VibrantLabs, 2024). And $R_{anc}$ is the LLM-as-Judge(Gu et al., 2025; Seed et al., 2025) score between the predicted and gold answers.
While $R_{anc}$ reflects the final objective, it is typically sparse or unstable in early training stages, making optimization difficult when used alone. To address this issue, we adopt an adaptive reward mixing strategy that enables a smooth curriculum-style transition from $R_{ans}$ to $R_{anc}$ during training:
At each update step $s$, the answer reward for a rollout $i$ is defined as a convex combination:
(6) $R^{t,i}_{a}\;=\;\alpha_{s}\,R_{ans}^{s,i}+(1-\alpha_{s})\,R_{anc}^{s,i},$
where $a_{s}\in(0,1)$ is a shared mixing weight for the entire batch.
To adaptively adjust $\alpha_{t}$, we track a batch-level estimate of model correctness. Specifically, let $R_{anc}^{t,i}\in[0,1]$ denote the raw correctness metric computed per rollout. We first compute the batch mean
(7) $\hat{c}_{s}\;=\;\frac{1}{bs}\sum_{i=1}^{bs}R_{anc}^{s,i},$
and maintain an exponential moving average (EMA)
(8) $\bar{c}_{s}\;=\;\beta\bar{c}_{s-1}+(1-\beta)\hat{c}_{s},\qquad\beta\in[0,1).$
The mixing weight is then defined via a monotone logistic gate:
(9) $\alpha_{s}\;=\;\sigma\!\big(\kappa(\tau-\bar{c}_{s})\big),$
where $\tau$ denotes a correctness threshold and $\kappa$ controls the sharpness of transition. As training progresses and $\bar{c}_{s}$ increases, $\alpha_{s}$ decreases smoothly, shifting the optimization focus from evidence consistency to answer correctness.
@@ 3.3.4. Overall Objective
(10) $R_{\mathrm{final}}=\begin{cases}\frac{R_{cite}+R_{pt}+R_{a}}{3},&\text{if format is valid,}\\ -1,&\text{otherwise.}\end{cases}$
@@
All components are jointly optimized using Group Relative Policy Optimization (GRPO)(Shao et al., 2024).
@@ 3.4. Startup Fine-Tuning
To initialize a reliable agent under our PoU reasoning framework, we generate expert trajectories by a strong LLM (GPT-5). These multi-step reasoning trajectories are reformatted into our standardized interaction schema, where each turn explicitly follows the PoU protocol.
@@
The resulting dataset thus consists of verifiable, well-grounded reasoning episodes that adhere to our explicit stepwise protocol. We fine-tune the base model on this curated corpus to cold-start the PoU agent.
@@ 4.2. Baselines
We compare PoU with representative deep-research agents covering retrieval–reasoning RAG, RL-based search, and multi-tool agent systems.
(1) Retrieval–Reasoning RAG.
Routing RAG dynamically selects retrievers (e.g., corpus, web, KG) based on query type and reasoning context, while All-in-One RAG fuses multiple retrievers into a unified interface for joint multi-source reasoning. Both follow the ReAct (Yao et al., 2023) framework with three reasoning–retrieval turns.
@@
Search-o1 (Li et al., 2025b), Search-R1 (Jin et al., 2025), and R1-Searcher (Song et al., 2025) perform multi-hop search via RL-guided sub-query generation, primarily over Wikipedia, followed by evidence summarization.
(3) Multi-Tool Agents.
DeepResearcher (Zheng et al., 2025) and Chain-of-Agent (CoA) (Li et al., 2025a) represent strong recent multi-tool agents operating over heterogeneous tool environments for deep research. CoA is trained by distilling trajectories from strong multi-agent systems and further refined with agentic reinforcement learning. DeepResearcher is an end-to-end reinforcement learning framework that trains LLM-based research agents in real-world web environments by learning to select and coordinate multiple web tools for search and browsing, using standard format and answer-correctness rewards, as in RL-based search agent baselines. For fair comparison, we train DeepResearcher from scratch using the same external tool set as PoU (denoted as DeepResearcher+), enabling evaluation under identical access to external knowledge sources.
(4) Distillation-Only PoU (SFT).
To further justify the necessity of explicit process-level RL supervision of PoU, we include an SFT-only baseline (PoU-SFT) trained on PoU-formatted expert trajectories, using the same model, tools, and data as PoU but without any RL rewards, thereby disentangling the effects of explicit process-level RL supervision from framework design and teacher–student distillation.
@@ 4.3. Implementation Details
We use GPT-4o-mini as the external judge model for evaluating answer–evidence alignment and factual consistency. All other agentic modules (browser, KG proxy, perturbations LLM) are implemented with Qwen-3-8B, deployed via vLLM inference servers. The training backbone model is Qwen-2.5-7B-Instruct. Prior to reinforcement learning, supervised fine-tuning is performed with a batch size = 256 for 3 epochs. For reinforcement learning, we adopt VERL with the GRPO algorithm. Each update uses 8 parallel rollouts, distributed across 32 × A100 GPUs. The $R_{pt}$ budget $B$ is 1. The maximum reasoning step is set to 10, and each retrieval call returns the top-5 documents or chunks per query. The KG proxy retrieves one-hop subgraph information from Wikidata and WikiPedia by ToG-2(Ma et al., 2025) system, providing semantically condensed neighborhood summaries as context. In the OOT experiments, we additionally introduce a local biochemical Proxy, which is built from a biochemical corpus derived from PubMed and serves as a local retrieval interface, serving as the fixed local knowledge base for domain-specific retrieval. The local_proxy is turn-off for non-medical domain tasks.
@@ 4.4. Main Results
Across all seven datasets—covering both in-domain (HotpotQA, 2Wiki) and out-of-domain (NQ, TQ, MuSiQue, Bamboogle, PopQA) benchmarks—our proposed Proof-of-Usefulness (PoU) agent consistently outperforms prior RAG and RL-based baselines. Notably, these gains extend beyond the training distribution: on all out-of-domain benchmarks, PoU achieves clear improvements in both factual F1 and LLM-as-judge (LM) metrics, demonstrating strong generalization ability.
Importantly, PoU also surpasses baselines on datasets such as NQ and TQ, which constitute in-domain data for the DeepResearcher baseline. This indicates that PoU’s improvements do not stem from memorizing domain-specific knowledge, but from stronger evidence–reasoning alignment. The model learns to effectively utilize retrieved evidence and to transfer its reasoning strategies to novel domains, capturing transferable reasoning competence rather than dataset-specific experience.
Results from DeepResearcher+ further show that naively expanding the action space by introducing additional information sources can degrade standard RL agents: they fail to exploit the added knowledge and suffer performance drops. This highlights the necessity of specialized training strategies for multi-source tool use. Moreover, PoU-SFT, despite being trained on LLM-guided expert trajectories, exhibits limited capability and degrades more severely on OOD tasks, suggesting that PoU’s gains do not arise from LLM-based rewards alone but from evidence-grounded training that teaches decision-making logic rather than simple imitation.
We also observe that All-in-One RAG consistently outperforms Routing RAG across all evaluation sets, corroborating findings from RAG in the Wild that naive or heuristic routing—without explicit optimization or supervision—is often inefficient or even counterproductive. While routing appears intuitively appealing by narrowing retrieval scope and reducing search cost, it risks prematurely discarding relevant information.
In our setting, the PoU agent operates in a multi-source environment where routing becomes an additional latent decision variable. Compared with traditional deep-research agents, this introduces a new layer of complexity: at each reasoning step, the agent must decide not only what to retrieve but also where to retrieve from. As the number of information sources grows, this expanded action space leads to inefficient exploration unless the routing policy is explicitly constrained. PoU addresses this challenge through structured reasoning and protocol-based training: by explicitly enforcing citation validity and reasoning consistency at every step, PoU regularizes the search process and prevents divergence in the enlarged action space. As a result, PoU can effectively leverage multi-source evidence without the instability and inefficiency characteristic of naive routing-based systems, yielding superior performance and generalization in multi-source environments.
@@ 4.5.1. Rewards Ablation
Figure 3 presents the ablation study of PoU under different reward configurations. We observe that increasing the perturbation budget to B=2B{=}2 leads to a more stable training trajectory and a higher final performance. This confirms the effectiveness of the perturbation-based reward $R_{pt}$: by actively perturbing the cited evidence, the model learns to rely on genuinely supportive information rather than superficially acknowledging tool responses. In addition, after removing adaptive reward mixing (i.e., using a static average of $R_{ans}$ and $R_{anc}$), compared to PoU, the learning curve shows a slower ascent in the early stage and a lower performance ceiling at convergence. This behavior reflects the stage-dependent role of reward signals. In the early phase, the model relies more heavily on $R_{\text{ans}}$ to accelerate learning by guiding coarse-grained process optimization. As training progresses and evidence–reasoning consistency is largely established, the model increasingly benefits from stronger answer-level supervision. This observation is well aligned with the original motivation of adaptive reward mixing. Furthermore, removing the answer–citation alignment term $R_{\text{ans}}$ leads to highly unstable training dynamics and frequent collapse in later stages. This suggests that explicitly linking the final answer with intermediate evidence provides an essential smoothing and grounding signal for policy optimization.
Figure 3. Ablation study of PoU. Comparison under different reward configurations: default perturbation budget B=1B{=}1, extended B=2B{=}2, and variants removing $R_{pt}$ or $R_{ans}$.
4.5.2. Out-of-domain-tool (OOT) and negative tool (NT) Experiments
To evaluate generalization under simultaneous domain and tool-set shifts, we conduct additional experiments in an out-of-domain (OOD) and out-of-domain-tool (OOT) setting. Specifically, we introduce tools that are never observed during training (a biomedical local search proxy (Local_Bio)). In addition, we introduce a negative tool (NT), which is deliberately task-irrelevant, to reflect realistic scenarios where certain tools are inapplicable and to assess the model’s robustness to interference as well as its ability to select appropriate tools. Experiments are conducted on two biomedical QA benchmarks, BioASQ (Krithara et al., 2023) and PQArefEval (bašaragin2024knowthatteachinggenerative).
Table 3. OOT evaluation on biomedical QA datasets (BioASQ, PQArefEval). “w/ NT” denotes the inclusion of an additional negative tool.
Dataset Zero-shot RAG DeepResearcher+ DeepResearcher DeepResearcher (w/ NT) PoU PoU (w/ NT)
BioASQ 13.5 25.8 37.1 38.9 31.8 45.6 43.9
PQArefEval 18.4 31.5 42.4 42.1 37.6 48.7 48.3
As shown in Table 3, DeepResearcher+, despite being trained with an expanded toolset, performs consistently worse than its original 2-tool counterpart (DeepResearcher) not only in OOD settings but also under the more challenging OOT regime across both biomedical datasets. Moreover, when an additional negative tool (NT) is introduced, the performance of DeepResearcher drops substantially, indicating a strong sensitivity to task-irrelevant tools. While PoU model achieves the best performance by a clear margin, demonstrating strong domain-level and tool-level generalization, while remaining robust even when the negative tool (NT) is introduced.
This result reveals an important phenomenon: as the number of available tools increases, the agent’s action space grows combinatorially. Without explicit process-level supervision, such an enlarged search space can easily lead to unstable policies or degenerate reasoning behaviors.
In contrast, through explicit evidence–reasoning-answer alignment training, the PoU training paradigm stabilizes learning under both domain and tool distribution shifts, substantially mitigating tool-call hacking. Crucially, PoU does not explicitly optimize for tool adaptation or robustness to task-irrelevant tools. Instead, by regulate a tight dependency between evidence usage, intermediate reasoning, and final answers, the training objective implicitly biases the trajectory toward tools that consistently support verifiable reasoning. As a result, robustness to unseen or task-irrelevant tools emerges as a extra bonus from PoU.